# Umberto Collodel Central Bank of Malta

### v1

@@ 4.1 Zero-Shot Prompt
The zero-shot specification provides the most direct test of whether LLM agents endowed only with fixed behavioral traits and no historical training context can reproduce the ebb and flow of market sentiment. Before turning to formal correlation measures, Figures 4(a) and 4(b) offer a visual narrative of the simulated output.
Panel 4(a) plots the share of agents predicting an increase, decrease, or no change in rates after each ECB press conference. Even in this naive setup, the agents’ directional calls broadly mirror well-known phases of ECB policy: a tightening bias before the Great Recession, a pronounced easing wave thereafter, a prolonged “on hold” stance at the short end during the zero lower bound (ZLB) years (contrasted with a slow grind lower for long maturities), followed by the sharp pandemic loosening and rapid tightening-partial reversal of the recent cycle. This ability to trace policy phases without explicit historical anchoring suggests that the agents’ built-in interpretive rules are sufficient to generate economically coherent directional patterns.
Panel 4(b) shows the cross-sectional standard deviation of forecasts. Dispersion contracts during periods of unambiguous policy stance, such as the post-crisis easing and much of the ZLB era, when consensus is high. It spikes at turning points—the onset of the Great Recession, the sovereign debt crisis, and recent tightening cycles—where disagreement arises less from direction and more from uncertainty over speed, magnitude, and endpoint. [^17] This is consistent with the ECB 2025 Monetary Policy Strategy Review (ECB,, 2025) findings. The report shows that, once forward guidance was abandoned, the sensitivity of OIS rates to macroeconomic news increased significantly, hence amplifying excess volatility. Relatedly, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge of aligning communication with policy moves. These patterns mirror survey-based and market-implied measures of disagreement, where even well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts if the terminal rate remains uncertain. [^18] See Gilchrist et al. (2014) and Bayer et al. (2022) for discussions on forecast disagreement and monetary policy uncertainty.
@@
Figure 4: Directional distribution and forecast dispersion from the zero-shot LLM simulation
(a) Percentage of forecast directions
(b) Standard deviation of forecasts
@@
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses ECB press conference transcripts and prompt 2.
@@
Figure 5(a) presents correlations with market measures. The 2-year tenor achieves the highest correlation (0.53), followed by the 10-year (0.44), and the 3-month (0.39). This hierarchy reflects the nature of information processed by the LLM and its impact on the yield curve. The 2-year tenor primarily reflects short- to medium-term policy expectations shaped directly by ECB forward guidance during press conferences. Unlike the 10-year yield, which incorporates long-term inflation expectations, structural growth prospects, and term premium fluctuations extending beyond ECB communication, [^19] See Kim and Wright (2005) for term premium dynamics. the 2-year segment is more narrowly focused on monetary policy expectations, resulting in a tighter linkage and higher correlation. The 3-month OIS rate exhibits low variance—especially during prolonged policy stasis such as the ZLB era—because well-anticipated policy leaves little room for market movement; thus, even accurate LLM decoding of policy tone cannot predict significant variation, constraining correlation.
@@
Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). Rolling correlation is calculated using a 12-month window.
For policymakers, a key consideration is not only the overall correlation level but also its stability and recent behavior, as shown in Figure 5(b). From 2006 to 2014, the 3-month tenor’s correlation was low and volatile, reflecting constrained short-term rate variation despite the LLM’s accurate interpretations. Starting in 2014, the ECB’s growing reliance on forward guidance is mirrored in steadily rising correlations at the 2-year tenor. The 10-year tenor shows more volatile but generally positive correlations, reflecting broader macroeconomic influences. Following 2022, correlations for both 2- and 10-year tenors surge amid aggressive ECB tightening to combat inflation, with the 2-year correlation consistently above 0.75. The decline in 10-year correlations in 2024 likely reflects markets looking beyond immediate tightening toward longer-term prospects.
Crucially, these results demonstrate that even a zero-shot LLM without historical training or calibration can effectively capture economically meaningful patterns within a behavioral-style simulation.
@@ 4.2 Few-Shot Prompt Results
The few-shot specification augments the prompt with the previous three ECB press conferences alongside their subsequent market-based measures of disagreement. This additional context is designed to help the LLM agents anchor their interpretations in patterns historically observed between ECB communication and subsequent market reactions, potentially enabling more informed directional and dispersion forecasts. Hansen et al., (2024) simulate economic forecasts of professional forecasters using LLMs, supplying them with detailed persona characteristics, real-time data, and past median Survey of Professional Forecasters projections. Inter alia, they find that removing past median forecasts largely decreases the prediction accuracy of the model, much more than removing precise human characteristics.202020”For some variables, the loss in performance is extreme. For example, the four-quarter ahead forecast of real GDP is 53% worse compared with the baseline, whereas the nowcast is worse by a factor greater than seven.” (Hansen et al.,, 2024). [^21] Their case can be thought of analogous to ours.
Figure 6: Directional distribution and forecast dispersion from the few-shot LLM simulation
(a) Percentage of forecast directions
(b) Standard deviation of forecasts
@@
Note: Each data point reflects the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty measure from Collodel and Kunzmann, (2025). Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses ECB press conference transcripts.
Panel 6 plots the share of agents predicting an increase, decrease, or no change in policy rates after each press conference. Relative to the zero-shot setup, the directional patterns remain broadly similar across major policy phases. Panel 6, instead, presents the cross-sectional standard deviation of forecasts. Here, the most visible difference from the zero-shot case is the large compression in dispersion, especially during periods of heightened policy uncertainty. [^22] For example, the upper bound of the distribution moves from 0.6 to 0.4 for the 10-years tenor. This implies that the historical examples nudge agents toward a more homogeneous reading of the ECB’s tone, aligning them more closely with recent precedents. While this reduces disagreement, it may also weaken the behavioral diversity that gives the simulation its interpretive richness.
Figure 7: Spearman correlation with market-based measures under the few-shot prompt
(a) Mean Spearman correlation by tenor
(b) Rolling Spearman correlation over time
Note: Correlation between the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of policy uncertainty from Collodel and Kunzmann, (2025). Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty Collodel and Kunzmann, (2025). Rolling correlation is calculated using a 12-month window.
Turning to the market correlations (Figure 7), the quantitative improvement is limited. The 2-year tenor sees a marginal decrease relative to the zero-shot baseline. The 10-year tenor, instead, moves in the opposite direction. At the very short end, the 3-month OIS correlation barely moves. All in all, these changes are not statistically significant.
The absence of improvement in the few-shot setting is consistent with two well-documented mechanisms in expectations formation. First, supplying a small and potentially unrepresentative set of recent examples can induce imitation and anchoring, compressing cross-sectional variance (e.g. Banerjee, (1992); Morris and Shin, (2002)). Second, when those examples originate from a different policy or macroeconomic regime, anchoring can miscalibrate forecasts, especially at turning points (Orphanides and Williams,, 2005). In our setup, a static few-shot context shared across agents, these forces plausibly reduce heterogeneity without yielding stronger market correlations.
Thus, that LLM agents exhibit stronger anchoring than market-based measures likely reflects the simulation design. The few-shot exemplars provide a fixed, shared reference point for all agents, with no access to broader or regime-specific information. LLMs in this setup cannot down-weight unrepresentative precedents, detect regime shifts, or benefit from incentives to differentiate. By contrast, human forecasters draw on heterogeneous information, adjust for changing regimes, and face strategic incentives that sustain disagreement. [^23] For example, price discovery actively amplifies minority views. For applications where distributional realism and market-like disagreement are more important than point accuracy, supplying a few recent examples can unintentionally induce consensus. Prompt design should therefore be tailored to each unique use case.
@@ 4.3 LLM-as-a-Judge
The Judge framework introduces an explicit feedback loop into the simulation. In this setting, a separate LLM evaluates each analyst’s past forecasting performance and rewrites the prompt to improve alignment with observed market reactions. This design allows the forecasting agents to adapt iteratively. We continue this back-and-forth for a total of 5 iterations, testing the new prompts on a training set, and then use every prompt produced for the test set. [^24] As for the other models, each prompt is run 5 times to get a robust standard deviation measure. This means we are running a total of 50 simulations: 5 simulations per iteration, 5 iterations, and 2 different datasets, train and test. This needs, however, also to be multiplied by the total number of ECB press conferences in our sample.
Figure 8 shows the average correlation across tenors for each judge refinement on both train and test. An analysis of the model’s performance reveals a high degree of robustness and effective generalization from the training to the test dataset. The Spearman correlation patterns are remarkably consistent across both sets, indicating that the model has learned fundamental relationships rather than simply memorizing the training data, thus avoiding significant overfitting. In both scenarios, performance peaks at specification 2, achieving a correlation of 0.58 in training and 0.57 in the test set, an economicaly significant increase compared to the naive prompt specification. Similarly, both datasets show a performance trough at specification 3. This consistency underscores the model’s reliability, while the decrease of performance underscores the importance of human supervision throughout the iteration process.
@@
The second prompt achieves the highest correlations in both training and test sets by providing structured guidance that links prediction dispersion directly to textual ambiguity (Table LABEL:tab:prompt_judge). According to the Judge, “its primary weakness is the lack of an explicit instruction linking the ambiguity or clarity of the ECB press conference transcript to the dispersion of the traders’ predictions.” [^25] In Collodel and Kunzmann, (2025), larger monetary surprises correspond to larger monetary uncertainty surprises. Interestingly, the model discovers this criterion independently. The model requires explicit cues such as deviations from market expectations, contrasting messages, or vague/hedging language. Subsequent prompts (3–5) show sharper degradation: prompt 3 enforces rigid dispersion calibration, while prompt 5 introduces a complex multi-factor uncertainty modeling. [^26] Prompt 4, if anything, returns closer to the initial specification. These conceptually sound instructions, however, constrain the model’s flexibility, resulting in overfitting on the training set or predictions that diverge from patterns observed in the test data.
Moving to a more detailed view, an analysis of the LLM-predicted standard deviation versus actual market volatility across different tenors and model iterations reveals a notable pattern (Figure 9). Focusing on fat-tailed events, such as the 2008–2009 financial crisis and the sovereign debt crisis, we observe that prompts 3–5 consistently fail to capture the full extent of market volatility. In contrast, prompts 1 and 2 tend to overshoot volatility, albeit with a certain lag (1-2 conferences). This behavior highlights that the model may struggle to accurately anticipate regime switches, particularly during episodes of extreme market stress, and independently from prompt engineering. [^27] Some small differences in the actual volatility measure across iterations is due to the LLM failing to return an output for a few conferences over all runs. Thus, those observations are removed from the plot.
@@
Note: Time series of LLM-predicted forecast dispersion against actual market-implied volatility across ECB press conferences, segmented by prompt iteration and tenor. Simulations use Gemini 2.5-Flash with a temperature of 1, and incorporate a Judge-based feedback loop that refines prompts over 5 iterations. Forecast dispersion is computed from 5 runs per prompt per conference. Market volatility is derived from OIS-based rate surprises. Red shaded areas represent the 90th percentile of the absolute difference between the two measures.

### v2

@@ 4.1 Zero-Shot Prompt
The zero-shot specification provides the most direct test of whether LLM agents endowed only with fixed behavioral traits and no historical training context can reproduce the ebb and flow of overall market sentiment. Before turning to formal correlation measures, Figures 4 and 5 offer a visual narrative of the simulated output.
@@
Figure 4: Percentage of forecast directions from the zero-shot LLM simulation
@@
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Forecast directions are aggregated across agents. The simulation uses a sample of 283 ECB press conference (June 1998-April 2025) transcripts and prompt 2.
Panel 4 plots the share of agents predicting an increase, decrease, or no change in rates after each ECB press conference. Even in this naive setup, the agents’ directional calls broadly mirror well-known phases of ECB policy: a tightening bias before the Great Recession, a pronounced easing wave thereafter, a prolonged “on hold” stance at the short end during the zero lower bound (ZLB) years (contrasted with a slow grind lower for long maturities), followed by the sharp pandemic loosening and rapid tightening-partial reversal of the recent cycle. This ability to trace policy phases without explicit historical anchoring suggests that the agents’ built-in interpretive rules are sufficient to generate economically coherent directional patterns.
Figure 5: Standard deviation of forecasts from the zero-shot LLM simulation
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2. Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
Panel 5 shows the cross-sectional standard deviation of forecasts, revealing distinct phases that align closely with monetary policy regimes and market stress episodes. Dispersion contracts systematically during periods of unambiguous policy stance, such as the post-crisis easing phase and much of the zero lower bound (ZLB) era, when forward guidance provided clear anchoring. In this period, the 3-month dispersion falls below 5 basis points, with longer tenors, instead, clustering around 10-15 basis points. Conversely, it spikes dramatically at policy turning points and during crisis episodes, touching sharp peaks during the 2008-2009 financial crisis (with the 2 and 10-year dispersion exceeding 40 basis points), the 2011-2012 sovereign debt crisis, and most recently during the 2022-2023 monetary policy tightening cycle, where disagreement reached unprecedented levels (roughly 60 basis points for the 2 and 10-year maturity). In the latter instance, disagreement arises less from directional confusion and more from uncertainty over policy speed, magnitude, and terminal endpoints. This result is consistent with the ECB 2025 Monetary Policy Strategy Review findings (ECB,, 2025). The report shows that, once forward guidance was abandoned, the sensitivity of OIS rates to macroeconomic news increased significantly, hence amplifying excess volatility. Moreover, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge for the ECB of aligning communication with policy moves. By contrast, following the introduction of the three-element reaction function framework in March 2023, the subsequent phase of the tightening cycle exhibits markedly subdued volatility to policy announcements. These patterns mirror survey-based and market-implied measures of disagreement, where even directionally well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts. [^17] See Gilchrist et al. (2014) and Bayer et al. (2022) for a thorough discussion on forecast disagreement and monetary policy uncertainty. [^18] At the end of 1998, volatility exhibits another peak for the 3-month tenor. In this period, the ECB avoided signaling its initial policy rate for the monetary union, deferring decisions to December. While acknowledging global risks from the Asian crisis, Russian default, and LTCM collapse, President Duisenberg reaffirmed the commitment to price stability, contrasting with the Fed’s three rate cuts between September and November. This ambiguity left markets uncertain and drove volatility in short-term instruments, mainly 1- to 3-month FRAs and DEM-based futures, since euro derivatives like Euribor futures and OIS did not yet exist (ECB,, 1999). The cross-tenor hierarchy remains remarkably consistent across all episodes, with 10-year uncertainty slightly larger than 2-year uncertainty, and both larger than the 3-month dispersion by 2-3 times, reflecting the increasing complexity of information processing and projection uncertainty at longer time horizons.
@@
Figure 6(a) presents correlations with market measures. The 2-year tenor achieves the highest correlation (0.53), followed by the 10-year (0.44), and the 3-month (0.39). This hierarchy reflects the differential nature of information processed by the LLM across the yield curve. The 2-year tenor primarily captures short- to medium-term policy expectations that are directly shaped by ECB communication during press conferences. In contrast, the 10-year yield incorporates broader factors extending beyond central bank communication, e.g., long-term inflation expectations, structural growth prospects, and term premium fluctuations, making it less tightly linked to the specific content of ECB press conferences. The 2-year segment’s narrower focus on monetary policy expectations, therefore, produces the strongest correlation with our LLM-based disagreement measure. At the short end, the 3-month OIS rate exhibits inherently low variance, particularly during prolonged policy stasis such as the ZLB era. Consequently, even when the LLM accurately interprets ECB communication tone, the constrained variation in 3-month rates limits the achievable correlation.
@@
Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2. Rolling correlation is calculated using a 12-month window.
For policymakers, a key consideration is not only the overall correlation level but also its stability and recent behavior, as shown in Figure 6(b). The rolling correlations reveal distinct regime-dependent patterns: during 2006-2014, correlations remained low and volatile across all tenors, reflecting constrained rate variation despite the LLM’s interpretive accuracy. The ECB’s adoption of forward guidance from 2014 onwards coincides with steadily rising correlations at the 2-year tenor, while 10-year correlations exhibit greater volatility due to broader macroeconomic influences beyond central bank communication. Most notably, the recent tightening cycle (2022-2024) produces a dramatic surge in performance, with 2-year correlations consistently exceeding 0.75 and 10-year correlations reaching their highest levels in the sample, demonstrating the framework’s particular effectiveness under current ECB communication practices. The subsequent decline in 10-year correlations during 2024 likely reflects markets’ forward-looking focus on longer-term economic prospects beyond the immediate policy cycle.
Crucially, these results demonstrate that even a simple zero-shot LLM approach, without historical training or fine-tuning, can capture economically meaningful patterns of market disagreement following central bank communication.
@@ 4.2 Few-Shot Prompt Results
We next test whether providing historical context improves performance by augmenting the prompt with the previous three ECB press conferences and their associated market-based disagreement measures. This few-shot approach aims to help agents anchor their interpretations in observed patterns between ECB communication and market reactions, following evidence that historical examples can improve LLM forecasting performance in other economic contexts (Hansen et al.,, 2024).
The few-shot specification does not materially improve correlation with market volatility compared to the naive model: correlations remain essentially unchanged across maturities, 2-year (0.52 vs. 0.48), 10-year (0.45 vs. 0.46), and 3-month (0.38 vs. 0.38), and none of these differences are statistically significant. The directional responses likewise mirror those produced under the zero-shot specification. Where the few-shot treatment matters is, however, in magnitude calibration. As Figure 7 shows, providing shared historical examples compresses forecast dispersion sharply, with maximum uncertainty for longer tenors falling from roughly 0.6 to about 0.3 percentage points, a level comparable to observed market-based volatility (Figure 1). This improved calibration of cross-sectional variance also makes the model’s first-moment forecasts much closer to reality.
Figure 7: Standard deviation of forecasts from the few-shot LLM simulation
@@
Note: Each data point reflects the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025), prompt 2 and provides the text of the three previous conferences as well as the ensuing market-based standard deviation of OIS rates. Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
Few-shot exemplars seem to help generate more realistic magnitudes for risk assessment and scenario analysis, yet they might also limit the dispersion that would otherwise capture market-like disagreement. In Banerjee, (1992), for example, agents confronted with common reference points tend to converge excessively on those anchors, leading to herding behavior even when the anchors are not fully representative of the true environment. By analogy, static exemplars in few-shot prompts may encourage the model to compress its distribution of forecasts around the precedent cases, improving calibration of magnitudes but reducing heterogeneity. At this stage, this remains more of an intuition than an empirical finding, and future work should test whether such a mechanism operates systematically and whether it can be mitigated.
@@ 4.3 LLM-as-a-Judge
The “Judge” framework introduces an explicit feedback loop into the simulation. In this setting, a separate LLM evaluates each analyst’s past forecasting performance and rewrites the prompt to improve alignment with observed market reactions. This design allows the forecasting agents to adapt iteratively. We continue this back-and-forth for a total of 5 iterations, testing the new prompts on a training set, and then use every prompt produced for the test set. [^19] As for the other models, each prompt is run 5 times to get a robust standard deviation measure. This means we are running a total of 50 simulations: 5 simulations per iteration, 5 iterations, and 2 different datasets, namely the training and test datasets. This, however, also needs to be multiplied by the total number of ECB press conferences in our sample. The stopping criterion emerged naturally: at the 6th iteration, the Judge attempted to reframe the task from predicting rate levels to basis point changes, which would have compromised our cross-sectional dispersion measurement framework. We therefore terminated optimization at iteration 5 to preserve methodological consistency.
Figure 8 reveals a striking pattern: the Judge’s refinements follow a clear performance arc across both training and test sets. Performance peaks dramatically at iteration 2, jumping from 0.47 to 0.58 in training, a substantial 11 percentage points improvement that holds remarkably well out-of-sample (0.57). This consistency across datasets suggests the Judge identified genuine signal rather than noise. Beyond iteration 2, however, the optimization process exhibits declining effectiveness. Iteration 3 marks a sharp deterioration, with correlations dropping to 0.31-0.37, before partially recovering in the 4th iteration (0.44-0.48) and then partially dropping again (0.36-0.35). This rollercoaster pattern reveals a key insight: more sophisticated prompts do not always yield better results. The Judge’s initial refinement successfully linked textual ambiguity to prediction dispersion, but subsequent ”improvements” appear to have over-engineered the task, constraining the model’s natural interpretive flexibility.
@@
The key to the Judge’s success in the first iteration lies in a deceptively simple insight. The second prompt achieves peak performance by establishing an explicit mapping between textual properties and prediction dispersion. As the Judge observes, the initial specification lacked ”explicit instruction linking the ambiguity or clarity of the ECB press conference transcript to the dispersion of the traders’ predictions.” (Table LABEL:tab:prompt_judge). Where the original prompt left this connection implicit, the refined version instructs agents that their collective disagreement ”must directly reflect the level of ambiguity, surprise, or conflicting signals within the ECB press conference transcript.” This finding echoes in part our earlier work showing that larger monetary surprises generate correspondingly larger uncertainty surprises (Collodel and Kunzmann,, 2025), though, remarkably, the Judge discovered this relationship independently through iterative refinement. The Judge’s diagnosis proves prescient: when given clear instructions to map vague language, contrasting messages, or deviations from expectations into wider prediction spreads, the LLM excels. However, subsequent iterations reveal the limits of automated optimization and demonstrate how optimization can become counterproductive. Prompt 3’s rigid dispersion calibration and Prompt 5’s elaborate multi-factor uncertainty framework, both conceptually sophisticated, actually constrain the model’s natural interpretive flexibility. The result is a classic case of over-engineering: theoretically sound but practically inferior performance that diverges from real-world patterns. Prompt 4, if anything, returns closer to the initial specification.
Moving to a more detailed view, an analysis of the LLM-predicted standard deviation versus actual market volatility across different tenors and model iterations reveals a notable pattern (Figure 9). The time series comparison exposes systematic differences in how prompt iterations respond to market stress, with consistent patterns emerging across the maturity structure.
Figure 9: Correlation with market-based measures across iterations - training set
@@
Note: Time series comparison of LLM-predicted forecast dispersion and actual market volatility following ECB press conferences, segmented by prompt iteration and tenor. Simulations employ Gemini 2.5-Flash with temperature = 1 and a Judge-based feedback loop that iteratively refines prompts over 5 iterations. Forecast dispersion represents the standard deviation across 5 simulation runs per prompt per conference. Market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). Red shaded areas indicate periods where prediction errors exceed the 90th percentile. Spearman correlation coefficients (r) between predicted and actual volatility are shown in red for each panel.
Focusing on fat-tailed events, such as the 2008–2009 financial crisis and the sovereign debt crisis, we observe distinct patterns across prompt iterations. Prompts 3–5 consistently fail to capture the full extent of market volatility, systematically undershooting across all three tenors. Despite their conceptual sophistication, these later iterations appear constrained by rigid calibration rules that limit their ability to generate extreme dispersion during crisis periods. Prompts 1 and 2 show the opposite behavior: they tend to overshoot volatility during these same periods, though with a 1-2 conference lag. Importantly, this overshooting preserves the maturity structure of actual market volatility; 10-year rates show the highest dispersion, followed by 2-year, then 3-month rates. This suggests that even when the model misjudges magnitude or timing, it captures the fundamental relationship between tenor length and uncertainty. The results highlight two key findings. First, the model struggles to anticipate regime switches during extreme market stress, regardless of prompt sophistication. Second, the LLM framework successfully internalizes basic term structure dynamics, maintaining realistic relationships across maturities even during periods of poor calibration.
To conclude, for central banks considering LLM-based communication analysis, these results offer both promise and caution. The technology can successfully model heterogeneous market reactions and even discover economic relationships independently. However, the optimization trajectory, from breakthrough to degradation within just a few iterations, underscores that human judgment remains indispensable in guiding AI systems toward practical applications.
5 Robustness
To validate our findings, we evaluate the robustness of the baseline LLM-based disagreement measure across three complementary dimensions that address key methodological concerns. First, we compare our approach with simple textual benchmarks to establish whether sophisticated agent-based modeling offers genuine advantages over basic complexity metrics. Second, we assess model architecture stability across different LLM systems (Gemini, ChatGPT, and Claude) to ensure our results reflect systematic patterns captured by this class of models rather than artifacts of a particular model’s design. Third, we examine prompt specification stability, addressing the well-documented fickleness of LLMs to minor wording changes.